# SWE-Synth: Synthesizing Verifiable Bug-Fix Data to Enable Large Language Models in Resolving Real-World Bugs

### v1

@@ 7.5.3 Impact of model size
[…]
@@ 7.5.4 Fix ground-truth extraction strategies
We investigate three strategies for extracting training patches from synthetic bug variants and evaluate their impact on APR performance. The approaches are described as follows: (1) Naive Reverse Patch Diff: the difference between the variant and the original selected component is used as the ground-truth fix (see Section 5). Then, we train a gold Retrieval Augmented Generation (gold RAG) model Jimenez et al. (2024) which provides the LLM model with the buggy file (gold) and its respective error log trace. (2) RAG with Rejection Sampling: Similar to the (1), this method uses candidate diffs; however, Qwen 2.5 Coder model is used to generate these candidates, then we select only those that successfully fix the bug. While this baseline removes unnatural fixes created by the reverse patch diff process by leveraging a LLM to produce the patches, it still lacks a detailed, intermediate repair steps. (3) Rollout Rejection Sampling (SWE-Synth): We derive the fix patch from a successful rollout trajectory of repair actions (as described in Section 5).
Quantitative Analysis. Table 9 summarizes the results. The naive reverse patch diff strategy achieves a very low empty patch rate (3.00%) but suffers from an extremely low resolve rate (0.67%). The RAG with rejection sampling one shows a modest improvement in resolve rate (3.00%), yet it still falls short of the zero-shot baseline of 11.00%. In contrast, the rollout rejection sampling strategy attains the highest resolve rate at 15.33%, showing the effectiveness of incorporating intermediate steps into the training dataset.
Qualitative Analysis. Figure 5 presents a variant generated from the pydantic project by reimplementing the __rich_repr__ method of the Representation class. The bug arises due to a type mismatch in the object yielded by the function. The respective test log and the generated intermediate steps (Steps 1–4) are shown. During the rollout, SWE-Synth does not know the bug location since it does not utilize the reverse patch diff. As observed, SWE-Synth employs the Moatless agent to analyze test logs and generate intermediate steps through reasoning. In Step 1, the agent attempts to retrieve relevant code context from the test logs. From Steps 2 to 4, it reasons about which components and how to modify them. Finally, it generates the patch, which differs from the original code.
@@ 8 Related Work
Second, manually curated datasets—such as SWE-Bench Jimenez et al. (2024) and SWE-Gym Pan et al. (2024b)—offer structured patches and executable environments with tests. These enable reasoning trace extraction but are inherently hard to scale due to the extensive human effort required.
Third, synthetic data generation offers a scalable alternative. Early efforts Xu et al. (2024); Luo et al. (2024); Wei et al. (2024); Yuan et al. (2025) have explored using LLMs to generate synthetic data, but these are often restricted to function-level tasks and lack full repository-level semantics, test-based verifiability, or repair traces. In contrast, our work focuses on creating repository-level APR datasets that are verifiable, context-rich, and scalable. Synthetic datasets have shown success in other domains such as healthcare Tang et al. (2023), mathematics Lu et al. (2024); Zhou et al. (2024a), and NLP Peng et al. (2023); Wang et al. (2023a); Abdin et al. (2024); Bartolo et al. (2021), but their application to executable SE tasks remains limited.
@@
Modern learning-based APR methods often frame bug fixing as a translation task using Neural Machine Translation (NMT) models Ye et al. (2023); Li et al. (2022a); Meng et al. (2023); Drain et al. (2021). While effective on simple benchmarks, they rely heavily on large-scale historical data and often suffer from generalization issues due to noise and lack of structured reasoning.
LLMs have introduced a paradigm shift, enabling zero-shot repair capabilities from natural language or code context Xia and Zhang (2022); Xia et al. (2023). This has inspired LLM-based agent frameworks for APR that simulate human workflows with structured tool usage and reasoning Yang et al. (2024); Örwall (2024); Antoniades et al. (2025); Zhang et al. (2024b); Xia et al. (2024b); Wang et al. (2024b); Phan et al. (2024); Nguyen et al. (2024). However, these agents typically rely on proprietary LLMs due to the lack of scalable, high-quality training data for open models—a gap our work directly addresses.
@@
SWE agent training requires executable and diverse environments. SWE-Bench Jimenez et al. (2023) provides gold patches but lacks execution scaffolding. R2E Jain et al. (2024) includes 246 instances with partial environment simulation. SWE-Gym Pan et al. (2024a) introduces executable tasks with human-curated test cases, enabling structured training but limiting scalability.
Our framework contributes a synthetic, verifiable, and scalable alternative. Inspired by synthetic dataset creation in other fields Long et al. (2024), our mutation and trace-generation pipeline offers rich repository-level training environments for agents.
Verifiers for Code Repair and Agent Decision-Making
Verifiers have become integral to improving patch quality. Execution-based verifiers Xia et al. (2024a) run generated patches against tests to validate correctness, while execution-free methods Pan et al. (2024a); Zhang et al. (2024a) score agent trajectories or use LLM-based judges. Each approach has strengths: execution-based methods offer precision but struggle with ambiguity; execution-free verifiers are flexible but may be misled by surface heuristics Gu et al. (2024). Recent work proposes hybrid strategies to combine the best of both Zhang et al. (2023); Ridnik et al. (2024).
We complement these efforts by generating data with built-in test signals and trace-level granularity, allowing verifiers to be trained and evaluated on both execution outcomes and reasoning paths.
@@ 9 Conclusion
SWE-Synth introduces a scalable approach using LLMs to generate synthetic bugs through code rewriting and rejection sampling. It also synthesizes intermediate repair steps, improving LLM-guided fixes. Evaluated on APR benchmarks (SWE-Bench and BugsInPy), an LLM trained on our dataset achieves comparable accuracy to manually collected datasets while having better scalability. An LLM agent trained on SWE-Synth outperforms SWE-Gym, achieving 15.33% resolve rate on SWE-Bench Lite compared to 13% by SWE-Gym.
@@
Data availability.
Our dataset/code are publicly available Anonymous (2025).

### v2

@@ 7.5.3 Impact of model size
$\%$ Resolved $(\uparrow)$ $\%$ Correct $(\uparrow)$ $\%$ Empty Patch $(\downarrow)$
@@ 7.5.4 Fix ground-truth extraction strategies
We investigate three strategies for extracting training patches from synthetic bug variants: (1) Naive Reverse Patch Diff: the diff between the variant and the original component is used as the ground-truth fix (Section 5), and a gold RAG model Jimenez et al. (2024) is trained with the buggy file and error log. (2) RAG with Rejection Sampling: Qwen 2.5 Coder generates candidate diffs, and only those that pass tests are retained, filtering out unnatural fixes but lacking intermediate repair steps. (3) Rollout Rejection Sampling (SWE-Synth): patches are derived from successful rollout trajectories of repair actions (Section 5).
Quantitative Analysis. Table 10 shows that reverse patch diff yields a very low resolve rate (0.67%) despite a low empty patch rate (3.00%). RAG with rejection sampling improves to 3.00% resolve rate but still underperforms the zero-shot baseline (11.00%). Rollout rejection sampling achieves the best performance with a 15.33% resolve rate, demonstrating the value of intermediate repair steps.
Qualitative Analysis. Figure 5 illustrates a pydantic bug involving the __rich_repr__ method. Without access to reverse diffs, SWE-Synth leverages the Moatless agent to analyze logs and iteratively generate repair steps. As shown, it retrieves context (Step 1), reasons about modifications (Steps 2–4), and ultimately produces a correct patch distinct from the original code.
@@ 8 Related Work
Second, manually curated datasets—such as SWE-Bench Jimenez et al. (2024) and SWE-Gym Pan et al. (2024b)—offer structured patches and executable environments with tests. These enable reasoning trace extraction but are inherently hard to scale due to the extensive human effort required. Recent efforts like SWE-Smith Yang et al. (2025) have explored automated pipelines for scaling data collection for SWE agents.
Third, synthetic data generation offers a scalable alternative. Early efforts Xu et al. (2024); Luo et al. (2024); Wei et al. (2024); Yuan et al. (2025) have explored using LLMs to generate synthetic data, but these are often restricted to function-level tasks and lack full repository-level semantics, test-based verifiability, or repair traces. Recent work on synthetic code datasets, such as KodCode Xu et al. (2025) for diverse coding challenges and HexaCoder Hajipour et al. (2024) for secure code generation, demonstrates the potential of oracle-guided synthetic training. However, their application to repository-level APR tasks with executable verification remains limited. Our work focuses on creating repository-level APR datasets that are verifiable, context-rich, and scalable, addressing this gap.
@@
Modern learning-based APR methods often frame bug fixing as a translation task using Neural Machine Translation (NMT) models Ye et al. (2023); Li et al. (2022a); Meng et al. (2023); Drain et al. (2021), building on advances in code understanding and representation learning Bui (2019); Nguyen et al. (2024b); To et al. (2024). While effective on simple benchmarks, they rely heavily on large-scale historical data and often suffer from generalization issues due to noise and lack of structured reasoning.
LLMs have introduced a paradigm shift, enabling zero-shot repair capabilities from natural language or code context Xia and Zhang (2022); Xia et al. (2023). This has inspired LLM-based agent frameworks for APR that simulate human workflows with structured tool usage and reasoning Yang et al. (2024b); Örwall (2024); Antoniades et al. (2025); Zhang et al. (2024a); Xia et al. (2024b); Wang et al. (2024a); Phan et al. (2024); Nguyen et al. (2024a). Notable approaches include RepairAgent Bouzenia et al. (2024), an autonomous LLM-based agent for program repair, and ThinkRepair Yin et al. (2024), which uses self-directed reasoning. RepairLLaMA Silva et al. (2025) demonstrates efficient fine-tuning strategies, while AlphaRepair Xia et al. (2024c) frames repair as code generation from bugs. Process-aware approaches like REPAIR Zhao et al. (2024) incorporate feedback during the repair process. However, these agents typically rely on proprietary LLMs due to the lack of scalable, high-quality training data for open models—a gap our work directly addresses.
Fault localization is a critical component of APR. Recent work has explored LLM-based fault localization Yang et al. (2024a), with AgentFL Qin et al. (2024) scaling these techniques to project-level context. Our synthetic dataset generation implicitly provides fault localization signals through test failures, enabling models to learn localization alongside repair.
@@
SWE agent training requires executable and diverse environments. SWE-Bench Jimenez et al. (2023) provides gold patches but lacks execution scaffolding. R2E Jain et al. (2024) includes 246 instances with partial environment simulation. SWE-Gym Pan et al. (2024a) introduces executable tasks with human-curated test cases, enabling structured training but limiting scalability. Our framework contributes a synthetic, verifiable, and scalable alternative. Inspired by synthetic dataset creation in other fields Long et al. (2024), our mutation and trace-generation pipeline offers rich repository-level training environments for agents.
Verifiers and Test Generation
Verifiers have become integral to improving patch quality. Execution-based verifiers Xia et al. (2024a) run generated patches against tests to validate correctness, while execution-free methods Pan et al. (2024a); Zhang et al. (2024b) score agent trajectories or use LLM-based judges. Each approach has strengths: execution-based methods offer precision but struggle with ambiguity; execution-free verifiers are flexible but may be misled by surface heuristics Gu et al. (2024). Recent work proposes hybrid strategies to combine the best of both Zhang et al. (2023); Ridnik et al. (2024).
Test generation has emerged as a complementary technique for APR verification. CodaMOSA Lemieux et al. (2023) combines search-based testing with LLMs to escape coverage plateaus, while Otter Ahmed et al. (2025) generates tests from issue descriptions to validate SWE patches. These approaches highlight the importance of test-based verification in APR workflows. We complement these efforts by generating data with built-in test signals and trace-level granularity, allowing verifiers to be trained and evaluated on both execution outcomes and reasoning paths.
@@ 9 Conclusion
In this work, we introduced SWE-Synth, a framework to overcome data scarcity in training LLMs for automated program repair. SWE-Synth employs a repository-level mutation strategy to generate large-scale, verifiable bug-fix data with human-like bugs, test cases for verification, and repair trajectories capturing the iterative debugging process.
Our contributions are threefold: (1) a scalable framework for producing vast synthetic bug data with minimal human oversight, (2) release of the SWE-Synth dataset containing bug-fix pairs, test suites, and intermediate repair steps, and (3) empirical validation showing an LLM agent trained on SWE-Synth achieves a 15.33% resolve rate on SWE-Bench Lite, surpassing the 13% rate of agents trained on manual datasets like SWE-Gym.
These results highlight synthetic data as a scalable, cost-effective alternative to manual collection, advancing autonomous software engineering and enabling more powerful, accessible automated bug-fixing solutions.